unseen:

Solve for x. x - 8 = -20 -28 -12 12 28

Comerpickles:

x=-12

iosangel:

fist add "8" to both sides because were doing the opposite of subtraction which is addition: x-8+8=-20+8=-12

iosangel:

She*

Comerpickles:

Oh god, I'm sorry, " she."

pokie:

i think the same thing

unseen:

thx

Comerpickles:

No problem and once again sorry about that Iosangel

iosangel:

Its ok XD